A car travels from Town A to Town B in 3hrs.
The same car travels 15 mph slower from Town B to Town A in 5hrs.
How fast was the car going to get from Town A to Town B?
I know know the Answer to the Question is 37.5 but I don't understand how you get the speed without knowing the distance.
: A car travels from Town A to Town B in 3hrs.
The same car travels 15 mph slower from Town B to Town A in 5hrs.
How fast was the car going to get from Town A to Town B?
I know know the Answer to the Question is 37.5 but I don't understand how you get the speed without knowing the distance.

Answer by scott8148(2483)

let x="speed from A to B", so x-15="speed from B to A"

d=rt __ the distance is the same in both directions

3(x)=5(x-15) __ 3x=5x-75 __ subtracting 5x __ -2x=-75 __ dividing by -2 __ x=37.5